6. George bought the following amounts of Stock A over the years: Stock A Stock A Stock A Date Purchased 11/21/1991 3/18/1997 5/22/2006 Number of Shares 1,000 500 750 Adjusted Basis $24,000 9,000 27,000 On October 12, 2017, he sold 1,200 of his shares of Stock A for $38 per share. a) How much gain/loss will George have to recognize if he uses the FIFO method of accounting for the shares sold? b) How much gain/loss will George have to recognize if he specifically identifies the shares to be sold by telling his broker to sell all 750 shares from the 5/22/2006 purchase and 450 shares from the 11/21/1991 purchase?
Answer #1

Answer a) Calculation of Gain in sale of ahare on FIFO basis

No. Of share sale = 1200 @38 per share

Cost of share = 1000 share @ 24 per share = $24000

= 200 share @ 18 per share = $ 3600

= $27600

Gain on Sale of share = 1200*38-27600 = 45600-27600 = $18000

Answer b)

Cost of Share = 750 share @36 = 27000

= 450 share @24 = 10800

= $37800

Gain on sale of share = 45600-37800

= $7800